OpenStudy (mendicant_bias):

Alright, now dealing with Surface Integral problems. Posting below momentarily.

OpenStudy (anonymous):

need help?

OpenStudy (mendicant_bias):

Yeah, just one moment.

OpenStudy (mendicant_bias):

G(x,y,z)=x, over the parabolic cylinder y = x^2, 0 leq x leq 2, 0 leq z leq 3

OpenStudy (mendicant_bias):

Alright, so I need to integrate the given function over the given surface. This is the first time I'm ever doing one of these problems, so it might be super easy, I just haven't tried yet until now; just posting it *immediately* in case I run into any misunderstandings, because I'm in a severe time crunch.

OpenStudy (mendicant_bias):

Alright, so G(x,y,z) is not defined parametrically, I need to get it into an either ecplicit, implicit, or parametrically defined form.

OpenStudy (anonymous):

yeah um i dont get this sorry

OpenStudy (mendicant_bias):

But yeah, reading, not yet sure how to advance/proceed from here.

OpenStudy (experimentx):

how do you find the surface area using integration?

OpenStudy (mendicant_bias):

Sort of learning/relearing how to do so, in better detail; Right now it looks like I have to take the magnitude of the cross product of two vectors to get something like a surface area differential?

OpenStudy (mendicant_bias):

http://i.imgur.com/MkNDPme.png

OpenStudy (experimentx):

yeah .. but that is after you parametrize the surface. There are couple of ways to do it.

OpenStudy (experimentx):

First thing is to find the region of integration on xy plane.

OpenStudy (mendicant_bias):

Alright, is this like the shadow method, or what are you talking about, exactly? My book has doing the vector cross product part first in its examples.

OpenStudy (experimentx):

look at the second and third method.

OpenStudy (experimentx):

that is more easier than you currently doing .. if you want to do it first way, write the parametric equation of paraboloic cylinder.

OpenStudy (mendicant_bias):

My book chooses to treat a problem with a near identical setup as my current problem as case 1 and solve it that way. Alright, then.

OpenStudy (anonymous):

@experimentX cam u help em

OpenStudy (mendicant_bias):

He is, joejoe.

OpenStudy (anonymous):

huh

OpenStudy (experimentx):

do you want to do it via parametric surface?

OpenStudy (mendicant_bias):

Yeah, to at the very least figure out/familiarize myself with the technique used in the first example. I'm getting how they took the cross product and stuff and how everything worked out, but confused about how to setup a parallel for this problem.

OpenStudy (mendicant_bias):

(?)

OpenStudy (mendicant_bias):

Alright, so parametrically defining a parabolic cylinder...I'm not sure how I should do that in this context, I could throw random tries at it, which gives me something like \[r = x, \ x^2 \ ,z\]

OpenStudy (experimentx):

do you know about parametric equations?

OpenStudy (mendicant_bias):

@satellite73

OpenStudy (mendicant_bias):

Any suggestions or anything? I'm not sure how to move forward with this yet.

OpenStudy (mendicant_bias):

Alright, then.. \[r =r(u,v)=ui+u^2j+vk\]

OpenStudy (mendicant_bias):

\[(u,v): [0,2], \ [0,3]\]

OpenStudy (mendicant_bias):

\[r_{u}, \ r_{v}=(1,\ 2u,\ 0)\ \ \ (0,\ 0,\ 1)\]

OpenStudy (experimentx):

can you imagine your surface .. or just draw rough picture of your surface?

OpenStudy (mendicant_bias):

Cross product between the two vectors gives you\[2u\ i-j \ ?\]

OpenStudy (mendicant_bias):

I dunno, I'm just going to keep doing what I'm doing, because I need to move fast on this, and doing what I'm doing seems to be getting closer to the answer quicker.

OpenStudy (mendicant_bias):

\[G(x,y,z)=x; \ \ \ G(r(u,v))=u\]

OpenStudy (mendicant_bias):

http://i.imgur.com/r7YOOpM.png

OpenStudy (mendicant_bias):

http://i.imgur.com/wp7Ez8L.png

OpenStudy (experimentx):

you don't need vector integral .. you only need the scalar surface integral.

OpenStudy (experimentx):

first thing you should do on multivariable calculus is ... visualize your equation in

OpenStudy (mendicant_bias):

Nope, apparently got it terribly wrong, but really want to get this problem done using *this* technique. My bounds of integration were correct, but my integrand was off.

OpenStudy (mendicant_bias):

I have no idea what your second to last post meant for me/how it alters the problem, can you point towards exactly what I did wrong? I need to figure out where that was.

OpenStudy (experimentx):

your surface is \( y = x^2 \) , \(0 \leq x \leq 2, 0 \leq z \leq 3 \)

OpenStudy (experimentx):

 this is how your surface looks like